OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4).

OpenStudy (anonymous):

You are looking for the linear equation based on those specifications. The equation of a lienar equation is \[y=mx+b\] m is the slope and b is the y intercept. \[y=,x+b\] \[-4=(-3*3)+b\] \[-4=-9+b\] \[-4+9=b\] \[5=b\] so now we get the y intercept value of b Then we plug that in to y=mx+b knowing m and b to get y=-3x+5

OpenStudy (anonymous):

Thank you(:

OpenStudy (anonymous):

No problem